OpenStudy (anonymous):

Solve the system of equations. x + 3y = −1 2x + 2y = 6

OpenStudy (sepeario):

Could you rearrange the first equation so we have an equation in terms of x (in the form of x=?)

OpenStudy (anonymous):

I am not sure

OpenStudy (sepeario):

Okay let me help you out then. If we move over the 3y onto the right hand side then we get \[x=-1-3y\]

OpenStudy (anonymous):

you can use substitution, and find x in x+3y=−1

OpenStudy (sepeario):

This form of solving a system of equations is known as substitution. Reason being, we can now substitute the value for x we have now into the second equation, hence eliminating the other variable.

OpenStudy (anonymous):

then when you find the x in that equation, substitute it for the x in 2x+2y=6

OpenStudy (anonymous):

to get y

OpenStudy (anonymous):

2x+3y = 1 2x = 1-3y x = (1-3y)/2 3(1-3y)/2 + 2y = -6 3-9y + 4y = -12 -5y = -15 y = 3 2x+3(3) = 1 2x + 9 = 1 2x = -8 x = -4 i hope this helps!